OpenStudy (anonymous):

WILL MEDAL AND FAN Becca has a biscuit recipe that uses 2/3 of a tablespoon of salt for every 3/4 of a cup of flour. In order to have enough biscuits for her family that is visiting, Becca is making a batch of biscuits using 3 cups of flour and ___ tablespoons of salt.

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} salt&flour \\\hline\\ \frac{2}{3}&\frac{3}{4}\\ x&3 \end{array}\implies \cfrac{\frac{2}{3}}{x}=\cfrac{\frac{3}{4}}{3}\impliedby \textit{solve for "x"} }\)

OpenStudy (studygurl14):

Helloo @Cira_Is_A_Neko ! Welcome to Openstudy! so, your question. For this, we can use proportions. 2/3 tbls of salt for every 3//4 cup of flour. We want to use 3 cups of flour, and don't know how much salt to use... \(\huge\frac{\frac{2}{3}}{\frac{3}{4}}=\frac{x}{3}\) Solve for x.

OpenStudy (anonymous):

@StudyGurl14 I think its x= 1/6

OpenStudy (anonymous):

Im not really good in math.

OpenStudy (studygurl14):

I'm sorry, but 1/6 is incorrect. Let's do the next step together. :)

OpenStudy (anonymous):

Oh, ok.

OpenStudy (studygurl14):

When you have proportions like this, you can do something called cross-multiply.   Do you see what I did?